OpenStudy (anonymous):

The CASH 6 lottery in Xanadu lets you pick six numbers in {1 . . 49}. How many ways are there for a player to get 4 or more number correctly.

OpenStudy (ash2326):

To win one need to get all six numbers correct. Do you agree with this?

OpenStudy (ash2326):

@nicole1234

OpenStudy (anonymous):

i suppose..but the question asks how many ways to get 4 correct...

OpenStudy (ash2326):

Yeah we want to get 4 or more So either we get 4 or 5 or 6 correct

OpenStudy (ash2326):

Do you know what's \[^aC_b?\]

OpenStudy (anonymous):

oh, that's right sorry. misread the problem

OpenStudy (anonymous):

this would be a combination with replacement problem, right?

OpenStudy (anonymous):

i'm not sure of the notation you used?

OpenStudy (ash2326):

\[^aC_b=\frac{a!}{b!(a-b)!}\] No of ways one can choose b items from a items

OpenStudy (ash2326):

Do you know this? It's not a replacement problem?

OpenStudy (anonymous):

ah, that's what i thought. Combination 'without' replacement is what i meant

OpenStudy (anonymous):

thanks for your help!

OpenStudy (ash2326):

Cool, Did you get it? You can solve it, can't you?

OpenStudy (anonymous):

i know how to find the number of ways to choose 6 items out of the 49, but what about the option of needing at least 4 correct?

OpenStudy (ash2326):

\[Total\ no.\ of\ ways=Ways\ to\ get\ 4\ correct+\ to\ get\ 5\ correct\ +\ to\ get\ 6\ correct\]

OpenStudy (anonymous):

Okay, so to double check, it would be the number of ways to choose 4 out of 49 plus the number of ways to choose 5 plus the number of ways to choose 6?

OpenStudy (ash2326):

Yeah correct:D

OpenStudy (ash2326):

No wait

OpenStudy (anonymous):

Thanks so much! I tend to get confused with word problems :)

OpenStudy (ash2326):

You have to choose like this, there are 6 correct and 43 incorrect To get 4 correct= ( choose 4 out of 6) \(\times\) ( choose the other two from the 43 incorrect) To get 5 correct= ( choose 5 out of 6) \(\times\) ( choose one from the 43 incorrect) To get 6 correct= ( choose 6 out of 6) \(\times\) ( choose none from the 43 incorrect)

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

i think so...i was just trying to work it out

OpenStudy (ash2326):

great, let me know if you get stuck. I'll help you in figuring out this:)

OpenStudy (anonymous):

so to find 4 correct I find (the number of ways to choose 4 out of 6)*(the number of ways to choose 2 out of 49) and then add that to the results of getting 5 correct and 6 correct?

OpenStudy (ash2326):

Yeah:) Whenever you have or it's equivalent to + \[or =>+\]

OpenStudy (anonymous):

makes sense :)

OpenStudy (ash2326):

@nicole1234 I forgot \[\large Welcome\ to\ Open\ Study:)\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

i'm getting 14707... does that sound about right?

OpenStudy (ash2326):

Let me check!!!

OpenStudy (ash2326):

6c4=15 43C2=903 6c5=6 43c1=43 6c6=1 43c0=1 \[15*903+6*43+1*1\] Did you get this?

OpenStudy (anonymous):

yes..but i accidently multiplied 903 by 16 instead of 15...ill fix it

OpenStudy (anonymous):

ok 13,804...

OpenStudy (ash2326):

Cool:D

OpenStudy (anonymous):

thank you so much!

OpenStudy (ash2326):

Glad to help:D:D